# Model Counting for Dependency Quantified Boolean Formulas

### v1

@@ 1 Introduction
While satisfiability is a central question in DQBF, many synthesis and verification tasks benefit from knowing how many solutions exist. Counting models can help debug and refine specifications: for instance, an unexpectedly large number of Skolem functions may suggest that the specification admits unintended behaviour. Model counters have been developed for QBF with one quantifier alternation (PlankMS24) as well as Boolean synthesis (kuldeep-skolem-24), and more recently, for general QBF (martina24).
@@
Note that functions in $\textup{EXP}$ may output doubly exponential numbers, which require exponentially many bits. Thus, the standard polynomial time Turing reductions for establishing $\textup{P}-$hardness, as in the case of #\#2-SAT in (Valiant79a; Valiant79b), is not appropriate for the class $\textup{EXP}$. To circumvent this issue, we introduce a new kind of polynomial-time reduction, called a poly-monious reduction (see Section 2 for the definition), which lies between classical parsimonious reductions and polynomial-time Turing reductions. Under poly-monious reductions, #\#2-SAT is still #\#P-complete (bannach_p_2025).
Another notion of hardness for counting problem requires the reduction to be parsimonious (Ladner89). However, since 2-SAT is NL-complete, #\#2-SAT is not #\#P-hard under parsimonious reduction, unless $\textup{NL}=\textup{NP}$. We believe that the notion of poly-monious reduction is well-suited for establishing $\textup{EXP}-$hardness, as it strikes a balance between parsimonious reductions and polynomial time Turing reductions in terms of strength.
@@
FOMC is often studied in data complexity setting, i.e., the FO sentence is fixed and the complexity is measured only in terms of the domain size. It is shown in (BeameBGS15) that there is a three-variable sentence such that the data complexity of its FOMC is $\textup{P}_{1}-$complete. For two-variable fragment, the data complexity drops to PTIME (toth-kuzelka-24; tim-kuzelka-23; BeameBGS15). The combined complexity is $\textup{P}-$complete, but assuming that the vocabulary is fixed (BeameBGS15). A tightly related problem to FOMC is query evaluation on probabilistic database, whose combined complexity is $\textup{P}-$complete (dalvi-suciu-04), but again, under the assumption of fixed vocabulary.
@@ 2 Preliminaries
Let $\Sigma$ be a finite alphabet. A poly-monious reduction from a function $F:\Sigma^{*}\to\mathbb{N}$ to another function $G:\Sigma^{*}\to\mathbb{N}$ is a polynomial time deterministic Turing machine $M$ together with a polynomial $p(s_{1},\ldots,s_{t})$ such that on input word $w$, $M$ outputs $t$ strings $v_{1},\ldots,v_{t}$ where $F(w)=p(G(v_{1}),\ldots,G(v_{t}))$.
@@ 3 Complexity of #\#DQBF
Before we can prove this, we need to introduce some further terminology. First, we recall the notion of succinct representation of graph introduced in (avi-succ-83). In such a representation, instead of being given the list of edges in a graph, we are given a boolean circuit $C(\bar{x},\bar{y})$, where $\bar{x},\bar{y}$ are vectors of Boolean variables of length $n$. The circuit $C$ represents a graph $G_{C}$ where the set of vertices is $\mathbb{B}^{n}$ and there is an edge oriented from $\bar{a}$ to $\bar{b}$, denoted $\bar{a}\to\bar{b}$, iff $C(\bar{a},\bar{b})=\top$.
@@
Finally, we need the notion of projection introduced in (skyum-valiant-V85). Intuitively, a projection is a special kind of polynomial time reduction where each bit $j$ in the output is determined either by the length of the input or by bit $i$ in the input where the index $i$ can be computed efficiently from index $j$ and the length of the input.
@@
(dqbf-sat23) Suppose there is a projection $\mathcal{A}$ that takes as input a CNF formula and outputs a graph. Then, there is a polynomial time algorithm that transforms a DQBF instance $\Psi$ to a circuit $C$ that succinctly represents the graph $\mathcal{A}(\exp(\Psi))$.
@@
$\beta\ :=\$$\bigwedge_{b,b^{\prime}\in\mathbb{B}}C(b,\bar{x},b^{\prime},\bar{x}^{\prime})\ \leftrightarrow\ (y_{1}^{b}\to y_{2}^{b^{\prime}})$
@@
We can also show that every $k$-DQBF can be reduced parsimoniously to uniform $k$-DQBF, which gives us the following corollary.
@@ 4 First-order Model Counting (FOMC)
It is implicit in (dqbf-fmcad22) that $\text{FOMC}_{\textsf{bin}}$ can be reduced to #\#DQBF and that the reduction is parsimonious. We will describe the idea here by an example. Consider the well known smoker-friend example for Markov Logic Networks (mln06):
$\Psi\ :=\$$\forall u\ \texttt{stress}(u)\to\texttt{smoke}(u)$
$\land\ \forall u\forall v\ \texttt{friend}(u,v)\land\texttt{smoke}(u)\to\texttt{smoke}(v)$
@@
The idea is to represent each of the predicates stress, smoke and friend with a Skolem function. We have $2n$ universal variables $\bar{x}_{1},\bar{x}_{2}$ in $\Phi_{n}$. The first block of $n$ variables $\bar{x}_{1}$ corresponds to $u$ and the second block $\bar{x}_{2}$ corresponds to $v$. It has $4$ existential variables, $y_{1},y_{2},y_{3},y_{4}$, corresponding to $4$ atoms $\texttt{stress}(u)$, $\texttt{smoke}(u)$, $\texttt{friend}(u,v)$ and $\texttt{smoke}(v)$. The dependency sets are $\bar{x}_{1}$, $\bar{x}_{1}$, $\bar{x}_{1}\cup\bar{x}_{2}$ and $\bar{x}_{2}$, respectively. The matrix of $\Phi_{n}$ is obtained by replacing each atom in $\Psi$ with its corresponding existential variable. Formally,
@@ 5 Algorithm for $\text{{2-DQBF}}$
High-level pseudocode is shown as Algorithm 1. First, using the reduction in (dqbf-fmcad24), we convert $\Phi$ to a transition system $(I,T)$, where $I$ is the formula for the initial states and $T$ is the formula for the transition relation, a brief summary of this transformation can be found in the supplementary material. From $(I,T)$, we can deduce whether $\Phi$ is satisfiable by constructing a formula $\varphi_{tr}$ that represents the transitive closure of $G_{\Phi}$ via BDD-based reachability. If it is not satisfiable, the algorithm immediately returns 0.
@@
Algorithm 1 Count the number of essential solutions for $\Phi$.
1:Transform $\Phi$ to a symbolic reachability instance $(I,T)$ using the transformation in  (dqbf-fmcad24).
@@
7:  Pick an arbitrary variable $X_{i,\bar{c}}$ from $R$.
8:  $C\leftarrow$ the connected component in $\widetilde{G}_{\Phi}$ that contain $X_{i,\bar{c}}$.
@@
10:       that respect the implications in $C$.
11:  Remove all the variables in $C$ from $R$.
12:  $N\leftarrow N\times N_{C}$.
13:return NN.
@@
into the transition relation $T$, for every $1\leq i\leq t$. We then check whether in the transition relation $T$ there is a cycle that contains contradicting literals. If there is such a cycle, we move to the next satisfying assignment of $A$ by “blocking” the assignment M in $A$. If there is no such a cycle, we extract the function $f$ for $y_{1}$ by employing the technique from (dqbf-fmcad24), add $f$ into $F$, and update $A$ by conjoining it with $C_{1}[\bar{z}_{1}/\bar{v}_{t+1}]$, where $\bar{v}_{t+1}$ are fresh variables, and $C_{1}$ is a Boolean formula extracted from $C$ that specifies only the literals associated with $y_{1}$.
@@
8:  $\triangleright$ Check if such assignments leads to no Skolem functions
@@
In Line 8 we force the assignment by adding into $T$ the edge $X_{1,\texttt{M}(\bar{v}_{i})}^{f^{(i)}(\texttt{M}(\bar{v}_{i}))}\to X_{1,\texttt{M}(\bar{v}_{i})}^{\neg f^{(i)}(\texttt{M}(\bar{v}_{i}))}$, for every $1\leq i\leq t$. In Line 9 we run the command reach from ABC to check whether there is a cycle containing contradicting literals. If there is such a cycle, the assignment M is blocked in Line 11. In Line 13 we extract a Skolem function using the technique from (dqbf-fmcad24). In Line 14 we count the number of solutions for the 1-DQBF instance after substituting $y_{1}$ with the new Skolem function $f$. Finally, in Line 16 we update the formula $A$ by conjoining it with the conjunction in Eq. (3).
@@ 6 Experiments
Figure 1: The two figures on the left shows the performance of the solvers on the PEC instances, the horizontal axis corresponds to the running time (s), and the vertical axis to the number of solved instances. The figure on the right shows the performance of the solvers on the 2-colorability instances, the horizontal axis corresponds to the number of bits of the graph in the instance, and the vertical axis to the running time (s).
@@
To evaluate our model counter, we generated a diverse family of benchmarks which we divide into three batches of instances.
  • PEC_opt: These are instances with dependency set size of 10 to 50. They are generated in a similar manner as in (dqbf-fmcad24). There are 370 instances in this batch. One third of these have 0 non-support variables.
  • PEC_small: These are instances generated from the ISCAS89 instances with dependency set size of 3 to 10 variables. There are 192 instances in this batch.
@@
Figure 1 shows that sharp2DQR fell short on PEC_small instances, but it is significantly better than Exp+ganak on PEC_opt instances. This is because the dependency set size is larger on PEC_opt instances, and since the expansion size is exponential to the dependency set size, our method, without expanding, performs better on larger instances. Most of the time spent by Exp+ganak is used on computing the expansion, and on many instances ganak finishes counting quite quickly after the expansion. sharp2DQR does not work well on small instances because sometimes BDD operations takes too long, and on the PEC_small instances, the number of unsatisfying models is small enough that enumeration is not too much of a problem. We also notice that for Exp+ganak, the performance of using cryptominisat5 and z3 is similar, but cryptominisat5 is better on the PEC_opt instances.
@@
More experimental results and analysis can be found in the supplementary material. We also compared both Exp+ganak and sharp2DQR against the latest FOMC tool WFOMC (wfomc_tool), where we encode counting the number of 2-coloring and independent sets on some specific graphs. In all instances, WFOMC can only handle model sizes of up to $4$, far lower than what sharp2DQR can handle, which in some instances is $2^{127}$. However, in the independent set counting instances, Exp+ganak performs better than sharp2DQR.
@@ 7 Concluding Remarks
Acknowledgement
This paper contains the supplementary material of the conference paper (us-aaai-26). We thank the reviewers of AAAI 2026 for their detailed and constructive comments on the initial version.
We would like to acknowledge the generous support of the Royal Society International Exchange Grant no. R3\233183, the National Science and Technology Council of Taiwan grant no. 111-2923-E-002-013-MY3, and the NTU Center of Data Intelligence: Technologies, Applications, and Systems grant no. NTU-113L900903.
More detailed proofs and experimental statistics
@@ Acknowledgement
In this supplementary material all logarithms have base $2$.

### v2

@@ 1 Introduction
While satisfiability is the central question in DQBF, many synthesis and verification tasks benefit from knowing how many solutions exist. Counting models can help debug and refine specifications: for instance, an unexpectedly large number of Skolem functions may suggest that the specification admits unintended behaviour. Model counters have been developed for QBF with one quantifier alternation (PlankMS24) as well as Boolean synthesis (kuldeep-skolem-24), and more recently, for general QBF (martina24).
@@
Note that functions in $\textup{EXP}$ may output doubly exponential numbers, which require exponentially many bits. Thus, the standard polynomial time Turing reductions for establishing $\textup{P}-$hardness, as in the case of #\#2-SAT in (Valiant79a; Valiant79b), are not appropriate for the class $\textup{EXP}$. To circumvent this issue, we introduce a new kind of polynomial-time reduction, called a poly-monious reduction (see Section 2 for the definition), which lies between classical parsimonious reductions and polynomial-time Turing reductions. Under poly-monious reductions, #\#2-SAT is still #\#P-complete (bannach_p_2025).
Another notion of hardness for the counting problem requires the reduction to be parsimonious (Ladner89). However, since 2-SAT is NL-complete, #\#2-SAT is not #\#P-hard under parsimonious reduction, unless $\textup{NL}=\textup{NP}$. We believe that the notion of poly-monious reduction is well-suited for establishing $\textup{EXP}-$hardness, as it strikes a balance between parsimonious reductions and polynomial-time Turing reductions in terms of strength.
@@
FOMC is often studied in the data complexity setting, i.e., the FO sentence is fixed and the complexity is measured only in terms of the domain size. It is shown in (BeameBGS15) that there is an $\text{FO}^{3}$ sentence such that the data complexity of its FOMC is $\textup{P}_{1}-$complete. For the two-variable fragment, the data complexity drops to PTIME (toth-kuzelka-24; tim-kuzelka-23; BeameBGS15). The combined complexity is $\textup{P}-$complete, but assuming that the vocabulary is fixed (BeameBGS15). A tightly related problem to FOMC is query evaluation on probabilistic databases, whose combined complexity is $\textup{P}-$complete (dalvi-suciu-04), but again, under the assumption of fixed vocabulary.
@@ 2 Preliminaries
Let $\Sigma$ be a finite alphabet. A poly-monious reduction from a function $F:\Sigma^{*}\to\mathbb{N}$ to another function $G:\Sigma^{*}\to\mathbb{N}$ is a polynomial-time deterministic Turing machine $M$ together with a polynomial $p(s_{1},\ldots,s_{t})$ such that on input word $w$, $M$ outputs $t$ strings $v_{1},\ldots,v_{t}$ where $F(w)=p(G(v_{1}),\ldots,G(v_{t}))$.
@@ 3 Complexity of #\#DQBF
Before we can prove this, we need to introduce some further terminology. First, we recall the notion of succinct representation of graphs introduced in (avi-succ-83). In such a representation, instead of being given the list of edges in a graph, we are given a Boolean circuit $C(\bar{x},\bar{y})$, where $\bar{x},\bar{y}$ are vectors of Boolean variables of length $n$. The circuit $C$ represents a graph $G_{C}$ where the set of vertices is $\mathbb{B}^{n}$ and there is an edge oriented from $\bar{a}$ to $\bar{b}$, denoted $\bar{a}\to\bar{b}$, iff $C(\bar{a},\bar{b})=\top$.
@@
Finally, we need the notion of projection introduced in (skyum-valiant-V85). Intuitively, a projection is a special kind of polynomial-time reduction where each bit $j$ in the output is determined either by the length of the input or by bit $i$ in the input, where the index $i$ can be computed efficiently from index $j$ and the length of the input.
@@
(dqbf-sat23) Suppose there is a projection $\mathcal{A}$ that takes as input a CNF formula and outputs a graph. Then, there is a polynomial-time algorithm that transforms a DQBF instance $\Psi$ to a circuit $C$ that succinctly represents the graph $\mathcal{A}(\exp(\Psi))$.
@@
$\beta\ :=\$$\bigwedge_{b,b^{\prime}\in\mathbb{B}}C(b,\bar{x},b^{\prime},\bar{x}^{\prime})\ \iff\ (y_{1}^{b}\to y_{2}^{b^{\prime}})$
@@
We can also show that every $k$-DQBF can be reduced parsimoniously to a uniform $k$-DQBF, which gives us the following corollary.
@@ 4 First-order Model Counting (FOMC)
It is implicit in (dqbf-fmcad22) that $\text{FOMC}_{\textsf{bin}}$ can be reduced to #\#DQBF and that the reduction is parsimonious. We will describe the idea here with an example. Consider the well-known smoker-friend example for Markov Logic Networks (mln06):
$\Psi\ :=\$$\forall u\forall v\ \ \texttt{stress}(u)\to\texttt{smoke}(u)$
$\qquad\land\ \texttt{friend}(u,v)\land\texttt{smoke}(u)\to\texttt{smoke}(v)$
@@
The idea is to represent each of the predicates stress, smoke, and friend with a Skolem function. We have $2n$ universal variables $\bar{x}_{1},\bar{x}_{2}$ in $\Phi_{n}$. The first block of $n$ variables $\bar{x}_{1}$ corresponds to $u$ and the second block $\bar{x}_{2}$ corresponds to $v$. It has $4$ existential variables, $y_{1},y_{2},y_{3},y_{4}$, corresponding to $4$ atoms $\texttt{stress}(u)$, $\texttt{smoke}(u)$, $\texttt{friend}(u,v)$ and $\texttt{smoke}(v)$. The dependency sets are $\bar{x}_{1}$, $\bar{x}_{1}$, $\bar{x}_{1}\cup\bar{x}_{2}$ and $\bar{x}_{2}$, respectively. The matrix of $\Phi_{n}$ is obtained by replacing each atom in $\Psi$ with its corresponding existential variable. Formally,
@@ 5 Algorithm for $\text{{2-DQBF}}$
High-level pseudocode is shown as Algorithm 1. First, using the reduction in (dqbf-fmcad24), we convert $\Phi$ to a transition system $(I,T)$, where $I$ is the formula for the initial states and $T$ is the formula for the transition relation. A brief summary of this transformation can be found in the appendix. From $(I,T)$, we can deduce whether $\Phi$ is satisfiable by constructing a formula $\varphi_{tr}$ that represents the transitive closure of $G_{\Phi}$ via BDD-based reachability. If it is not satisfiable, the algorithm immediately returns 0.
@@
Algorithm 1 Count the number of essential solutions for $\Phi$
1:Transform $\Phi$ to a symbolic reachability instance $(I,T)$ using the transformation in  (dqbf-fmcad24)
@@
7:  Pick an arbitrary variable $X_{i,\bar{c}}$ from $R$
8:  $C\leftarrow$ the connected component in $\widetilde{G}_{\Phi}$ that contain $X_{i,\bar{c}}$
@@
10:       that respect the implications in $C$
11:  Remove all the variables in $C$ from $R$
12:  $N\leftarrow N\times N_{C}$
13:return NN
@@
into the transition relation $T$, for every $1\leq i\leq t$. We then check whether in the transition relation $T$ there is a cycle that contains contradicting literals. If there is such a cycle, we move to the next satisfying assignment of $A$ by “blocking” the assignment M in $A$. If there is no such cycle, we extract the function $f$ for $y_{1}$ by employing the technique from (dqbf-fmcad24), add $f$ into $F$, and update $A$ by conjoining it with $C_{1}[\bar{z}_{1}/\bar{v}_{t+1}]$, where $\bar{v}_{t+1}$ are fresh variables, and $C_{1}$ is a Boolean formula extracted from $C$ that specifies only the literals associated with $y_{1}$.
@@
8:  $\triangleright$ Check if such assignments lead to no Skolem functions
@@
In Line 8, we force the assignment by adding into $T$ the edge $X_{1,\texttt{M}(\bar{v}_{i})}^{f^{(i)}(\texttt{M}(\bar{v}_{i}))}\to X_{1,\texttt{M}(\bar{v}_{i})}^{\neg f^{(i)}(\texttt{M}(\bar{v}_{i}))}$, for every $1\leq i\leq t$. In Line 9 we run the command reach from ABC to check whether there is a cycle containing contradicting literals. If there is such a cycle, the assignment M is blocked in Line 11. In Line 13 we extract a Skolem function using the technique from (dqbf-fmcad24). In Line 14, we count the number of solutions for the 1-DQBF instance after substituting $y_{1}$ with the new Skolem function $f$. Finally, in Line 16 we update the formula $A$ by conjoining it with the conjunction in Eq. (3).
@@ 6 Experiments
Figure 1: The two figures on the left show the performance of the solvers on the PEC instances, the horizontal axis corresponds to the running time (s), and the vertical axis to the number of solved instances. The figure on the right shows the performance of the solvers on the 2-colorability instances; the horizontal axis corresponds to the number of bits of the graph in the instance, and the vertical axis to the running time (s).
@@
To evaluate our model counter, we generated a diverse family of benchmarks, which we divided into three batches of instances.
  • PEC_opt: These are instances with a dependency set size of 10 to 50. They are generated in a similar manner as in (dqbf-fmcad24). There are 370 instances in this batch. One third of these have 0 non-support variables.
  • PEC_small: These are instances generated from the ISCAS89 instances with a dependency set size of 3 to 10 variables. There are 192 instances in this batch.
@@
Figure 1 shows that sharp2DQR fell short on PEC_small instances, but it is significantly better than Exp+ganak on PEC_opt instances. This is because the dependency set size is larger on PEC_opt instances, and since the expansion size is exponential to the dependency set size, our method, without expanding, performs better on larger instances. Most of the time spent by Exp+ganak is used on computing the expansion, and on many instances ganak finishes counting quite quickly after the expansion. sharp2DQR does not work well on small instances because sometimes BDD operations take too long, and on the PEC_small instances, the number of unsatisfying models is small enough that enumeration is not too much of a problem. We also notice that for Exp+ganak, the performance of using cryptominisat5 and z3 is similar, but cryptominisat5 is better on the PEC_opt instances.
@@
More experimental results and analysis can be found in the appendix. We also compared both Exp+ganak and sharp2DQR against the latest FOMC tool WFOMC (wfomc_tool), where we encode counting the number of 2-colorings and independent sets on some specific graphs. In all instances, WFOMC can only handle model sizes of up to $4$, far lower than what sharp2DQR can handle, which in some instances is $2^{127}$. However, in the independent set counting instances, Exp+ganak performs better than sharp2DQR.
@@ 7 Concluding Remarks
Acknowledgements
This paper is the extended version of the conference paper with same title (us-aaai-26). We thank the AAAI’26 reviewers for their detailed and constructive comments on the initial version.
We would like to acknowledge the generous support of the Royal Society International Exchange Grant no. R3\233183, the National Science and Technology Council of Taiwan grant no. 111-2923-E-002-013-MY3 and 114-2221-E-002-183-MY3, and the NTU Center of Data Intelligence: Technologies, Applications, and Systems grant no. NTU-113L900903.
Appendix: More detailed proofs and experimental results
@@ Acknowledgements
In this appendix we present some of the missing proofs in the main body. We also present more experimental results and the statistics. All logarithms have base $2$.
@@ Appendix I More experimental results
Some statistics.
Most of the solved instances contains one component, the largest number of components that sharp2DQR solved is 244. For PEC_opt, the largest count we solved is around $2^{1.8\text{e}19}$, while the median is $2^{9220}$. For solutions over support variables, a lot of them contains one partial solution per component, and a few of them contains $2$ to $2^{589824}$ total partial solutions over the support variables. For PEC_small, the largest count we solved is $2^{1088}$, while the median is $2^{4}$. For solutions over support variables, again, a lot of them contain one partial solution per component, and a few of them contain $2$ to $2^{514}$ total partial solutions over the support variables.